If log(a+c), log(a+b), log(b+c) are in AP and 1/a, 1/b, 1/c are also in AP, then find the value of a+b given a,b,c>0

Grade:9

1 Answers

Arun
5 years ago
Dear Aryan
 
2 log (a+ b) = log (b+ c) + log (c + a)
 
(a+ b)^2 = (b+ c) (c+ a)
 
also
 
2/ b = 1/a + 1/ c
 
b = 2 ac/(a+c)
 
a+ c = 2 ac /b
 
now you can find the value